### v1

# Yintong Huo‡§, Michael R. Lyu

Abstract
Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in automated front-end engineering, e.g., generating UI code from visual designs. However, existing front-end UI code generation benchmarks have the following limitations: (1) While framework-based development becomes predominant in modern front-end programming, current benchmarks fail to incorporate mainstream development frameworks. (2) Existing evaluations focus solely on the UI code generation task, whereas practical UI development involves several iterations, including refining editing, and repairing issues. (3) Current benchmarks employ unidimensional evaluation, lacking investigation into influencing factors like task difficulty, input context variations, and in-depth code-level analysis.
To bridge these gaps, we introduce DesignBench, a multi-framework, multi-task evaluation benchmark for assessing MLLMs’ capabilities in automated front-end engineering. DesignBench encompasses three widely-used UI frameworks (React, Vue, and Angular) alongside vanilla HTML/CSS, and evaluates on three essential front-end tasks (generation, edit, and repair) in real-world development workflows. DesignBench contains 900 webpage samples spanning over 11 topics, 9 edit types, and 6 issue categories, enabling detailed analysis of MLLM performance across multiple dimensions. Our systematic evaluation reveals critical insights into MLLMs’ framework-specific limitations, task-related bottlenecks, and performance variations under different conditions, providing guidance for future research in automated front-end development. Our code and data are available at https://github.com/WebPAI/DesignBench.
Index Terms:
Multimodal Large Language Models, Code Generation, Web Development.
I Introduction
Converting webpage designs into functional UI code is a critical yet labor-intensive step in web development. MLLMs have demonstrated remarkable performance on visually rich code generation tasks [1, 2, 3, 4, 5], creating new opportunities for automated design-to-code conversion that replicates webpage elements, layouts, text, and colors.
Several benchmarks have been proposed to evaluate MLLMs on front-end code generation. These benchmarks typically involve either synthesizing webpage code via LLMs (e.g., WebSight [6] and Web2Code [7]), or curating code by cleaning real-world webpages (e.g., Webcode2M [8] and Design2Code [9]). However, while these evaluation benchmarks are effective in measuring certain MLLM capabilities, they do not adequately represent the complex challenges faced by developers in real-world development scenarios.
(1) Lack of front-end framework integration. Current benchmarks fail to incorporate front-end frameworks, such as React, Vue, and Angular, which are integral to modern web development (as indicated in 2025 trends [10]). Consequently, MLLM capabilities for practical framework-based website development remain unexplored.
(2) Insufficient task coverage. Existing benchmarks inadequately cover the comprehensive spectrum of front-end development tasks. As illustrated in Fig. 1, development encompasses not only initial code generation from UI designs, but also iterative code editing for design refinement and repairing design-related issues [11, 12]. However, current evaluations focus exclusively on the generation phase, neglecting the critical edit and repair phases.
Figure 1: Pipeline of MLLM-based automated front-end engineering.
(3) Limited evaluation dimensions. Most existing studies assess the overall quality of the generated webpage, lacking a detailed, multi-dimensional analysis of MLLM performance. Such an analysis could consider varying difficulty levels, input contexts, and code-level attributes like correctness and reusability. The comprehensive evaluation is crucial for identifying strengths and weaknesses of MLLMs, thereby fostering more reliable models for practical deployment.
To address these limitations, we propose DesignBench, a comprehensive multi-framework multi-task benchmark for evaluating MLLMs across multiple dimensions. DesignBenchcovers three popular front-end frameworks (React, Vue, Angular) and encompasses three key tasks: design generation, edit, and repair. The benchmark contains 900 webpage samples spanning more than 11 topics, 9 edit types, and 6 issue categories, enabling detailed analysis of MLLM performance across difficulty levels, context, and code dimensions.
Our extensive experiments yield several key findings: (1) Framework-specific limitations. MLLMs exhibit substantially lower performance in framework-based development compared to vanilla HTML/CSS. They struggle with framework-specific syntax, such as JSX parsing (React), template syntax (Vue), and TypeScript architecture (Angular). Critically, these models fail to effectively leverage framework features and component-based paradigms, resulting in diminished code reusability. (2) Task-specific bottlenecks. Distinct performance bottlenecks were identified across the three core development tasks. Design generation task suffers from visual rendering inaccuracies and compilation errors, while design editing and repair tasks are primarily constrained by shortcomings in code localization and UI issue identification capabilities. (3) Performance variability under different conditions. MLLM efficacy significantly degrades under more challenging conditions, including large UI images for generation, complex instructions for editing, and severe UI issues for repair. Notably, an input context analysis on design edit and repair tasks reveals that code-only inputs consistently outperform image-only inputs. Multimodal combinations provide minimal additional improvement, suggesting that code representations convey more precise semantic information for these modification tasks than visual inputs.
TABLE I: Comparison of existing UI-to-Code benchmark and DesignBench. Vanilla refers to plain HTML/CSS.
Benchmark Sample Source Framework Task Dimension
Pix2code [13] 1742 Synthetic Vanilla Generation
WebSight [6] 823K Synthetic Vanilla Generation
Web2Code [14] 60K Synthetic Vanilla Generation
WebCode2M [8] 20K Real-world Vanilla Generation
Design2Code [9] 484 Real-world Vanilla Generation
DesignBench (Ours) 900 Real-world Vanilla, React, Vue, Angular Generation, Edit, Repair Difficulty, Context, Code
Our main contributions are summarized as follows:
  • We introduce the first comprehensive multi-framework, multi-task benchmark for evaluating MLLMs’ capabilities in automated front-end engineering across HTML/CSS, React, Vue, and Angular frameworks.
  • We conduct an extensive evaluation of nine leading MLLMs across three fundamental tasks. The analysis considers multiple dimensions, including task difficulty (visual complexity, instruction complexity, and issue severity), input context modalities (code-only, image-only, and multimodal), and code metrics (correctness and reusability).
  • We reveal key insights into MLLMs’ framework-specific limitations, identify task-dependent performance bottlenecks, and characterize performance variations under diverse conditions. These results provide important guidance for future research and practice in front-end development.
  • To facilitate further research, we publicly release the code and data on Github [15].
II Background
II-A Web Development Process
A typical web application front-end development process includes the following stages:
  1. 1.
    Design stage: designers create high-fidelity mock-ups using prototyping tools such as Sketch [16] and Axure [17] during design stage.
  2. 2.
    Development stage: this phase involves transforming the design concepts into a functional application through coding. The development stage typically consists of the implementation of the GUI and the underlying functionalities. As shown in Fig. 1, the developer first prompts the MLLM to generate the front-end code based on the UI-Mockup. However, the version one code did not fully comply with the UI-Mockup. Then the user asked MLLM to edit the code to change the background color and add buttons to generate version two. Then during the actual deployment process, the front-end engineer found a display issue on the UI, that is, a collision between the avatar and the name, and asked MLLM to fix it and generate version three code.
II-B Front-end Web Development Framework
Front-end frameworks are collections of pre-written, reusable code that provide a foundation for building the user interface of a website. These frameworks can save time and improve the quality of the final product by providing a standardized set of components and templates for front-end development. In 2025, React, Vue, and Angular have emerged as the predominant frameworks in web development [10], with 39.5%, 15.4%, and 17.1% popularity [18], respectively.
III Related Work
III-A UI Code Generation
UI code generation techniques fall into three categories: Deep Learning (DL) based methods, Computer Vision (CV) based methods, and Multimodal Large Language Model (MLLM) based methods. (1) DL-based methods: [19, 20, 21, 22, 23] leverages CNNs to automatically prototype software GUIs. Pix2code [13] utilizes CNNs and LSTM to extract features from GUI images to generate a domain-specific language (DSL). [24] implements an encoder-decoder framework with an attention mechanism to generate the DSL. (2) CV-based methods: Sketch2Code [25] inputs hand-drawn sketches into object detection models to learn the object representation, which is read by the UI parser to generate code for targeted platforms. REMAUI [26] identifies user interface elements via optical character recognition (OCR) techniques and then infers a suitable user interface hierarchy and exports the results as source code. (3) MLLM-based methods [27, 28, 29, 30]: to solve the element omission distortion and misarrangement problems during UI code generation, DCGen [28] proposes a divide-and-conquer-based approach to generate the code of the submodules separately and then assemble them to construct the full webpage based on MLLMs. DeclarUI [29] applies the element segmentation method to accurately generate elements and page transition graphs to prompt MLLMs to generate mobile app UI with jump logic. While prior works achieve decent performance in UI-to-code, none of them address the conventional framework-based development and other essential front-end procedures like design edit and repair.
III-B Benchmarks for UI Code Generation
Many benchmarks [31, 32, 33, 34] have been proposed to evaluate the code generation, understanding and reasoning capabilities of LLMs, but there are only a limited number of benchmarks for UI code. To improve MLLMs’ UI-to-Code capabilities, Pix2code introduced a domain-specific language (DSL) and leveraged CNN and LSTM architectures to translate UI mockups into DSL code. However, due to the inherent limitations of DSL approaches, Pix2code suffers from poor scalability and limited real-world applicability. WebSight advanced the field by synthesizing high-quality HTML code for training, while Web2Code [7] proposed the Webpage Code Generation Benchmark (WCGB) to systematically evaluate MLLMs’ HTML parsing capabilities. Despite these contributions, both benchmarks rely on synthetic data, which may not capture the complexity and variability of real-world web development scenarios. Design2Code [9] addressed this limitation by manually curating 484 authentic web pages from the Common Crawl dataset, constructing the first real-world benchmark for design-to-code evaluation. Building upon this foundation, WebCode2M [8] significantly expanded the scale with 20,000 samples, providing both comprehensive training data for model development and robust test sets for evaluation. Other Benchmarks like Interaction2Code [35] and MRWeb [36] mainly focus on the interactive and multi-page web application generation, which is out of our scope.
Table I presents a comparison between existing benchmarks and our, in terms of the size, collection method, framework, target tasks, and evaluation dimension. DesignBench is distinct in incorporating varied front-end frameworks, multiple tasks, and diverse evaluation dimensions.
IV DesignBench
IV-A Task Definition
As shown in Fig. 1, the automated front-end development mainly contains the following three tasks: “Design Generation”, “Design Edit” and “Design Repair”.
Design Generation ($\mathcal{T}_{G}$). The objective of design generation is to generate expected code based on the UI Mockups. Formally, given a UI design image $I$, the task aims to generate corresponding UI code $C$ such that $\mathcal{T}_{G}:I\rightarrow C$. The input contains the UI design image $I$, and the output is the UI code $C$ that accurately reproduces the visual layout and styling.
Design Edit ($\mathcal{T}_{E}$). The goal of the design edit is to generate front-end code that complies with user modification instructions. Given the original UI design image $I_{o}$, original UI code $C_{o}$, and user instruction $T$ described in natural language, the task produces modified code $C_{new}$ such that $\mathcal{T}_{E}:(I_{o},C_{o},T)\rightarrow C_{new}$. The input contains the original UI design image $I_{o}$, original UI code $C_{o}$, and user instruction $T$, while the output is the updated code $C_{new}$ incorporating the requested modifications.
Design Repair ($\mathcal{T}_{R}$). The goal of the design repair is to repair the UI code with display issues. Given the problematic UI code $C_{p}$, the problematic UI image $I_{p}$, the task generates repaired UI code $C_{r}$ such that $\mathcal{T}_{R}:(C_{p},I_{p})\rightarrow C_{r}$. The input contains the problematic UI code $C_{p}$ and image $I_{p}$, the output is the repaired code $C_{r}$ that resolves visual design issues.
IV-B Data Collection
Figure 2: Pipeline of DesignBench  construction.
Design Generation. For websites developed by frameworks, we collect samples from GitHub and the top 500 globally visited websites. Existing reports [10, 18] show that the three most popular front-end development frameworks are React, Vue, and Angular, so we collect data for these three frameworks to build our benchmark. (1) Github projects. We search for “React projects”, “Vue projects” and “Angular projects” to get a summary list of web projects with different frameworks, then we identify 152 popular projects with deployed links and higher star counts. These projects represent various real-world website uses, ranging from commercial product front-end websites to blogs, with 4055 average star counts and an average of 996 forks. Their popularity has proven their usefulness and quality. (2) Top 500 globally visited websites. These websites are ranked by Moz [37], and we collect 158 webpages that adopt frameworks as our dataset.
For websites without applying frameworks, we sample 120 webpages with different lengths from the webcode2m dataset [8]. Ultimately, we compile a dataset consisting of 430 webpages. After getting the link of the webpage, we apply single-file tool [38] to save the webpage as a stand-alone file, then replace the images in the web page with placeholders, and finally use selenium [39] to take screenshots for constructing the input design image $I$ for task $\mathcal{T}_{G}$.
Design Edit. To obtain real-world user instructions for UI design modification, we crawl 541 React-based projects from Vercel’s V0 platform [40] and 1,349 Vue-based projects from Vue0 [41]. These projects contain comprehensive interaction histories, including user instructions and corresponding UI code implementations.
We first filter the projects that can compile successfully with at least two iteration rounds. For each project with consecutive editing steps $v_{1}$ and $v_{2}$, we construct evaluation pairs where the input consists of the previous version’s code $C_{v1}$, rendered UI image $I_{v1}$, and the user instruction $T_{v1\rightarrow v2}$ that describes the desired modifications. The ground truth comprises the updated code $C_{v2}$ and its corresponding rendered UI image $I_{v2}$ that implements the requested changes.
However, some edit steps feature ambiguous user instructions or poorly UI modifications that fail to satisfy the intended requirements. To filter out low-quality samples, we employ five PhD students with three years of front-end development experience to conduct a comprehensive assessment.
Specifically, the annotators classify user instructions into three levels of clarity: clear, moderate, and ambiguous, based on detailed assessment of instruction specificity, actionability, and comprehensibility. Similarly, they evaluate the quality of modified UI and categorize them into terrible, good, and excellent based on how well the changes align with the given instructions. The detailed annotation guidelines are in [15]. The final classifications for both instruction clarity and UI modification effectiveness are determined through a majority voting process among the annotators.
Finally, we curate 359 high-quality React and Vue samples that receive clear instruction ratings and excellent modification quality scores. To obtain Angular and vanilla HTML/CSS samples, we randomly select 146 samples from this curated set and translate them into Angular and vanilla HTML/CSS code implementations using GPT-4o [42]. Following the automated translation, the same five PhD students verify the correctness of the translated implementations, and make necessary modifications to ensure the code meet users’ instructions.
Design Repair. After collecting the webpages of the above two tasks, the five PhD students screen out the UI samples with design issues. They follow comprehensive guidelines to identify display problems, including: (1) Layout issues such as misaligned elements and incorrect positioning; (2) Visual inconsistencies including improper spacing, incorrect font sizes, or color scheme violations; (3) Component rendering problems such as missing elements, overlapping content, or distorted images. The detailed annotation guidelines are in [15].
Finally, we get 111 webpages with problematic UI code $C_{p}$ and image $I_{p}$. To obtain the repaired code $C_{r}$ for evaluation, 111 webpages are evenly assigned to the five PhD students to manually fix the UI display issues. Each student spent about 5 hours to complete the repair of all assigned samples.
IV-C Data Annotation
Figure 3: The UI issue types in DesignBench. The red bounding box marks the issue location.
Design Generation. The annotators are employed to annotate the topics of the webpages based on their functions.
Design Edit. The annotators are instructed to annotate both the type and difficulty level of UI modifications. The modification type encompasses two dimensions: the operation type and the corresponding UI attribute type being adjusted. The operation types include three fundamental categories: Add (introducing new UI elements), Change (modifying elements), and Delete (removing elements). The UI attribute types subject to adjustment comprise six main categories: text (including content, font, and typography modifications), color (encompassing background colors, text colors, and accent colors), position (spatial arrangement and layout adjustments), size (dimensional scaling and resizing operations), shape (geometric modifications and structural changes), and component-level (holistic modifications affecting entire UI components). The difficulty level reflects several factors, primarily the number of UI elements that require adjustment, the complexity of the interdependencies between modified elements, and the scope of cascading changes required to maintain design consistency and functionality throughout the interface.
Design Repair. The annotators are employed to annotate the UI issues. Aftering collecting the webpages with display issues, we first randomly select 25% samples for analysis and then discuss, revise, and refine the UI issue type until everyone reaches a consensus. During annotating new data, if encountering a new issue type, annotators will communicate and update issue type in time to guide subsequent annotations. Finally, we classify 6 types of UI display issues as follows:
  • Occlusion. Elements are hidden or partially covered by other elements, making content inaccessible or invisible to users. This includes overlapping components, modal dialogs blocking content, or elements positioned behind others. As shown in Fig. 3(a). The “Doctor Name” box partially covers the portrait frame.
  • Crowding. Too many elements are packed into a small space without adequate spacing, making the interface feel cluttered and difficult to navigate. Fig. 3(b) shows an example that the “Start Test” and “Reset” buttons are tightly packed together.
  • Text overlap. Text content overlaps with other text or UI elements, making it unreadable or causing visual confusion. Fig. 3(c) shows that the “AI Chat” text and “Support Bot” text are overlapped.
  • Alignment. Elements are not properly aligned with each other or the overall layout grid, creating a disorganized appearance. As shown in Fig. 3(d), the “Feature 1” and “Feature 2” titles are not aligned with icons.
  • Color and contrast. Poor color choices that affect readability or accessibility, including insufficient contrast between text and background, or color combinations that are difficult for users with visual impairments to distinguish. As illustrated in Fig. 3(e), the text color and background color are too close to each other, making it difficult to distinguish.
  • Overflow. Content extends beyond its intended container boundaries, causing horizontal scrollbars, cut-off text, or elements appearing outside their designated areas. For example, the text “Don’t have an account” exceeds the login container.
IV-D Benchmark Statistics
Table II presents the sample counts of the three tasks with three frameworks, the Vanilla denotes the webpage developed by vanilla HTML/CSS.
TABLE II: Sample counts of three frameworks on three tasks.
Framework Design Generation DesignEdit Design Repair
React 109 108 28
Vue 118 105 27
Angular 83 66 28
Vanilla 120 80 28
Total 430 359 111
(a) Topic Distribution.
(b) Issue Distribution.
Figure 4: Category distribution of Design Generation and Design Repair task.
(a) Operation Distribution.
(b) Visual Distribution.
Figure 5: Edit type distribution of Design Edit task.
Fig. 4(a) shows that DesignBench covers a diverse range of web topics with more than 11 types, including information, homepage, tool, product, news, and so on. This extensive topic coverage demonstrates that DesignBench encompasses a broad range of aspects of digital interface design, ensuring comprehensive evaluation across different domain requirements.
Fig. 5 shows the distribution of the edit type of design edit tasks. The operation distribution reveals three main categories: Change (55.8%), Add (38.1%), and Delete (6.1%). The corresponding visual modifications encompass diverse types, including component-level changes (27.9%), text modifications (17.1%), position adjustments (16.0%), color alterations (15.2%), shape modifications (13.4%), and size adjustments (10.4%), indicating that design edit tasks cover a broad spectrum of modification requirements.
Fig. 4(b) shows the distribution of the UI issues of design repair task. The issue types span multiple dimensions with alignment being most prevalent (42.2%), followed by crowding (18.7%), occlusion (18.1%), overflow (11.4%), and other visual defects. This wide range of issue types demonstrates that design repair tasks address comprehensive quality assurance across multiple design issues dimensions.
V Experiment Setup
V-A Models
The studied MLLMs are listed in Table III. We select six state-of-the-art LLMs that have been widely explored in multimodal tasks, three from open-source models, namely Pixtral [43], Qwen [44], and LLama [45], three from commercial models like Gemini [46], GPT [42] and Claude [47]. In configuring the MLLM models, we set the temperature to 0 and the maximum number of tokens output as the upper limit of MLLMs’ maximum output token. All other parameters are kept at their default settings. The detailed prompts of the three tasks are available in our code repository [15]. The entire benchmark evaluation incurs an average API cost of \$52 per model on average. The average processing times per sample with a single thread, are 49 seconds for generation, 29 seconds for editing, and 25 seconds for repair, respectively.
V-B Metric
We evaluate the performance of the model on DesignBench from three types of metrics:
Visual Metrics. CLIP [48] is applied to measure the semantic similarity between the generated and original webpages.
Code Metrics. (1) Compilation Success Rate (CSR). This metric represents the percentage of generated code that compiles successfully without errors. Assume that the total number of samples is N and the number of samples compiled successfully is S, then $CSR=\frac{S}{N}$. (2) Code Modification Similarity (CMS). We employ the Jaccard similarity [49] to quantify the precision of code modifications on design edit and design repair tasks by comparing the sets of modified line numbers between the ground truth and generated code. Let $A$ represent the set of line numbers modified in the ground truth code and $B$ represent the set of line numbers modified in the generated code. The CMS is formally defined as: $CMS(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
MLLM-as-Judge Metrics. MLLMs have shown great performance in assisting judges across diverse modalities [50, 51]. Therefore, we prompt GPT-4o [42] to determine whether the model meets the user’s requirements on the design edit task and resolve the design issues on the design repair task, and output an MLLM score between 0 and 10 with detailed explanations (0-3 denotes the poor edit/repair, 4-6 denotes partial edit/repair, 7-8 denotes Good eidt/repair and 9-10 denotes excellent edit/repair). For design edit task, we sample 359 samples and validate this MLLM score through human evaluation, which can achieve an average accuracy of 95.54%. For design repair task, we sample 111 samples and validate this MLLM score through human evaluation, which can achieve an average accuracy of 91.89%.
TABLE III: Studied Multimodal Large Language Models.
Base Model Model Abbreviation Size
Pixtral [43] Pixtral-12B-2409 Pixtral-12B 12B
Pixtral [43] Pixtral-large-latest Pixtral-124B 124B
Qwen [44] Qwen2.5-VL-7B-Instruct Qwen-7B 7B
Qwen [44] Qwen2.5-VL-72B-Instruct Qwen-72B 72B
LLama [52] Llama-3.2-11B-Vision-Instruct Llama-11B 11B
LLama [52] Llama-3.2-90B-Vision-Instruct LLama-90B 90B
Gemini [46] Gemini-2.0-Flash Gemini-2.0 -
GPT [42] GPT-4o-2024-11-20 GPT-4o -
Claude [47] Claude-3-7-sonnet-20250219 Claude-3.7 -
V-C Research Questions
  • RQ1: (Performance across tasks) How do MLLMs perform across distinct front-end tasks?
  • RQ2: (Performance across frameworks) What is the comparative performance of MLLMs when applied to different development frameworks?
  • RQ3: (Influence of difficulty) How does varying task difficulty impact MLLM performance?
  • RQ4: (Influence of context) To what extent do different input contexts affect MLLMs’ performance?
  • RQ5: (Limitation analysis) What are the primary limitations of MLLMs in developing framework-based webpages?
VI Experiment Results
TABLE IV: The model performance on DesignBench under different tasks and different frameworks. Bold numbers on a dark red background indicate the maximum values, and an underline with a light red background denotes the second-best value.
Metric Framework Claude GPT Gemini Llama Pixtral Qwen
Claude-3.7 GPT-4o Gemini-2.0 Llama-90B Llama-11B Pixtral-124B Pixtral-12B Qwen-72B Qwen-7B
Design Generation
CLIP React 0.8083 0.7637 0.7611 0.7040 0.6401 0.7395 0.6168 0.7790 0.0875
Vue 0.8319 0.7734 0.6897 0.5323 0.3243 0.7811 0.7434 0.6836 0.0452
Angular 0.6024 0.5963 0.6006 0.5327 0.4891 0.6324 0.4876 0.5149 0.0851
Vanilla 0.8132 0.7683 0.7588 0.6404 0.6304 0.7403 0.7043 0.7597 0.7411
Compilation React 0.9541 0.9725 0.9083 0.9450 0.8991 0.9725 0.8532 0.9541 0.1284
Vue 0.9746 0.9492 0.8390 0.7458 0.4915 0.9746 0.9407 0.8559 0.0678
Angular 0.6867 0.7108 0.7108 0.7349 0.6988 0.7590 0.6024 0.6265 0.1205
Design Edit
MLLM Score React 8.1759 8.0093 7.8148 6.1574 4.8148 8.0185 7.6111 8.0833 1.8796
Vue 8.3619 8.1810 8.0571 6.2571 3.1333 8.0381 7.0190 7.5714 2.2952
Angular 8.0152 8.2879 9.1364 5.6515 5.1212 8.6818 7.8030 8.1970 2.0152
Vanilla 9.1500 9.2250 9.0250 7.7000 6.5750 9.1125 8.6000 9.1250 5.9125
CMS Vue 0.4050 0.3698 0.3276 0.2104 0.0655 0.3046 0.2394 0.3276 0.0862
React 0.4659 0.5246 0.3710 0.2637 0.1819 0.4093 0.4120 0.4398 0.0815
Angular 0.6829 0.6099 0.6392 0.4700 0.3621 0.3264 0.2867 0.6018 0.1367
Vanilla 0.3439 0.3394 0.2905 0.1946 0.1582 0.3651 0.2770 0.3209 0.1635
Compilation React 1.0000 0.9815 1.0000 0.9167 0.7963 0.9907 1.0000 0.9907 0.4815
Vue 0.9810 0.9429 0.9524 0.9143 0.5905 0.9619 0.9048 0.9333 0.4286
Angular 0.9091 0.9091 1.0000 0.8636 0.7727 0.9848 0.9242 0.9091 0.3333
Design Repair
MLLM Score React 6.7857 6.3571 6.3214 4.2143 2.7500 6.4643 5.3571 5.6429 0.8929
Vue 6.5926 6.2593 6.0741 4.7778 3.5185 6.3704 6.0370 6.0370 0.4815
Angular 6.8571 5.9286 5.2857 4.6429 3.2500 6.5357 5.6429 6.5000 0.0000
Vanilla 7.1786 7.0714 7.3214 5.7143 5.7857 6.9643 6.6786 6.8929 3.8571
CMS React 0.4827 0.2752 0.1755 0.0448 0.0473 0.2272 0.0905 0.1866 0.0417
Vue 0.3065 0.2524 0.1782 0.0501 0.0474 0.2230 0.1557 0.1131 0.0127
Angular 0.5719 0.5073 0.3968 0.3099 0.2688 0.2618 0.2546 0.5563 0.0000
Vanilla 0.2287 0.1637 0.1630 0.0365 0.0690 0.1398 0.1188 0.1446 0.0277
Compilation React 1.0000 1.0000 1.0000 0.9286 0.9643 1.0000 0.9643 0.9286 0.2857
Vue 1.0000 1.0000 0.9630 1.0000 0.8889 1.0000 1.0000 1.0000 0.1111
Angular 0.9286 1.0000 1.0000 0.7857 0.8571 1.0000 1.0000 0.9286 0.0357
VI-A RQ1: How do MLLMs perform across distinct front-end tasks?
Table IV presents the performance of nine MLLMs on three tasks in three front-end frameworks React, Vue, and Angular. Vanilla denotes the webpage developed by vanilla HTML/CSS.
Among all, Claude-3.7, GPT-4o, Gemini-2.0, and Pixtral-124B are the top-performing MLLMs in three tasks. Claude-3.7 achieves the highest performance across most metrics, including superior CLIP scores (0.6024-0.8319) for Design Generation and exceptional MLLM scores for Design Edit (8.01-9.15) and Design Repair (6.59-7.17). GPT-4o follows closely with strong CLIP scores (0.5963-0.7734) and outstanding compilation rates (0.7108-0.9725), alongside competitive MLLM scores across all tasks. Gemini-2.0 demonstrates solid performance with CLIP scores ranging from 0.6006-0.7611, reliable compilation success rates consistently above 0.71, and strong MLLM scores in Design Edit (7.81-9.13) and Design Repair (5.28-7.32). Pixtral-124B rounds out the top tier with competitive performance across multiple metrics, achieving strong CLIP scores (0.6324-0.7811), excellent compilation rates (0.7590-0.9746), and robust MLLM scores for Design Edit (8.01-9.11) and Design Repair (6.37-6.96).
Finding 1: Among the evaluated MLLMs, Claude-3.7, GPT-4o, Gemini-2.0, and Pixtral-124B consistently demonstrated top-tier performance across the three tasks.
Larger variants consistently outperform their smaller counterparts within the same family. This is evident in comparisons such as Llama-90B versus Llama-11B, Pixtral-124B versus Pixtral-12B, and Qwen-72B versus Qwen-7B comparisons. The performance advantages are particularly pronounced in complex tasks requiring code localization and visual understanding, suggesting that increased model capacity enhances essential web development capabilities.
Finding 2: Larger models consistently outperform smaller variants, demonstrating that increased model capacity enhances web development capabilities.
In Design Generation tasks, MLLMs face two primary bottlenecks: compilation errors and visual inaccuracies. The compilation rates reveal significant framework-dependent challenges, with Angular showing the lowest success rates (0.6867-0.7590 for top models) compared to React and Vue ($>$0.83 for top models). Additionally, the moderate CLIP scores (0.5963-0.6324 even for best performers) indicate substantial room for improvement in generating visually accurate webpage layouts. The performance gap between vanilla HTML (highest CLIP scores) and framework-based implementations further suggests that the syntactic complexity in modern frameworks exacerbates challenges in both compilation and visual rendering.
In Design Edit and Design Repair tasks, the primary bottleneck of MLLMs is the accurate localization of code segments requiring modification. This is evidenced by CMS scores that are lower than compilation rates across all models. In Design Edit, even top-performing models like Claude-3.7 achieve CMS scores of only 0.3439-0.6829, despite maintaining compilation rates above 0.9. Similarly, in Design Repair, CMS scores range from 0.2287-0.5719 for Claude-3.7. These results highlight a substantial difficulty in localizing target code, even when the generated code successfully compiles.
Finding 3: MLLMs exhibit task-specific bottlenecks: Design Generation is challenged by compilation errors and visual inaccuracies, while Design Edit and Repair are mainly limited by deficiencies in code localization.
VI-B RQ2: What is the comparative performance of MLLMs when applied to different development frameworks?
(a) CLIP.
(b) Compilation.
Figure 6: Performance of webpages implemented by different frameworks. y-axis denotes the framework used to actually implement the webpage and x-axis represents the framework used by the model.
TABLE V: Performance under different difficulty levels.
Model Design Generation Design Edit Design Repair
Easy Medium Hard Easy Medium Hard Easy Medium Hard
Claude-3.7 0.83 0.76 0.73 8.64 8.32 7.61 7.21 6.81 6.93
GPT-4o 0.79 0.74 0.69 8.73 7.95 8.19 6.93 6.46 5.53
Gemini-2.0 0.80 0.73 0.70 9.05 7.84 7.51 7.07 6.21 4.19
Llama-90B 0.68 0.64 0.62 6.54 6.83 5.68 5.35 4.90 4.39
Llama-11B 0.65 0.60 0.59 5.88 5.05 5.24 3.91 4.07 2.41
Pixtral-124B 0.78 0.72 0.71 8.69 7.91 6.98 7.30 6.56 5.66
Pixtral-12B 0.70 0.66 0.68 8.26 7.63 6.71 6.76 5.84 5.30
Qwen-72B 0.76 0.69 0.69 7.78 6.81 6.72 7.12 6.06 5.57
Qwen-7B 0.55 0.53 0.52 3.46 2.83 2.71 1.09 1.63 1.05
To further explore the model’s proficiency across different front-end frameworks, we evaluate MLLMs’ ability to implement webpages using various framework combinations.
Fig. 6 presents the average CLIP scores and compilation success rates of nine MLLMs across different framework combinations. The results reveal distinct performance patterns across frameworks. MLLMs consistently achieve optimal performance when implementing webpages using vanilla HTML/CSS, attaining the highest CLIP scores above 0.72 and perfect compilation success rates. In contrast, Angular-based implementations demonstrate the poorest performance, with compilation success rates ranging from 0.6-0.7 and CLIP scores between 0.45-0.55. React and Vue frameworks show intermediate performance levels, with both achieving reasonable compilation rates and moderate CLIP scores, though still inferior to vanilla implementations.
Finding 4: MLLMs demonstrate the strongest performance with vanilla HTML/CSS, followed by React and Vue, but exhibit significant challenges with Angular implementations.
VI-C RQ3: How does varying task difficulty impact MLLM performance?
We categorize samples into different difficulty levels to systematically evaluate MLLMs’ performance across varying complexity scenarios. The difficulty assessment criteria differ for each task to reflect their distinct challenges.
For Design Generation task, which require generating code from scratch, difficulty is primarily determined by the visual complexity of webpages. We classify webpages smaller than 1000$\times$1000 pixels as easy, those larger than 2000$\times$2000 pixels as hard, and intermediate sizes as medium difficulty.
In Design Edit task, where models must modify existing code according to user instructions, difficulty correlates with the complexity and scope of the modification instructions. We adopt the annotator-provided difficulty labels from Section IV-C as our ground truth difficulty classification.
For Design Repair task, difficulty is assessed based on the severity of UI issues, quantified by the extent of code modifications required to resolve the identified problems. Tasks requiring modifications to more than 30 lines of code are classified as hard, those requiring fewer than 10 lines as easy, and intermediate cases as medium difficulty.
Table V shows the CLIP score for Design Generation task and MLLM score for Design Edit and Design Repair task under different difficulty levels.
Table V reveals distinct difficulty-related performance patterns across the three design tasks. In Design Generation, where difficulty stems from image complexity and size, top models show moderate degradation from  0.79-0.83 (Easy) to  0.69-0.73 (Hard). Design Edit tasks, with difficulty determined by the number of required operations, exhibit more pronounced drops, particularly for complex multi-operation scenarios where top models decline from 8.64-9.05 to 7.51-8.19. Design Repair shows the most severe degradation, where UI issue severity and required code modifications cause performance to plummet from 6.93-7.21 (Easy) to 4.19-6.93 (Hard), with smaller models experiencing catastrophic failures. These patterns indicate that visual complexity moderately affects generation, operational complexity significantly impacts editing, while code-level debugging presents the steepest performance barriers for current MLLMs.
Finding 5: MLLM performance degrades when confronted with large UI images in Design Generation, complex instructions in Design Edit, and severe UI issues in Design Repair tasks.
VI-D RQ4: To what extent do different input contexts affect MLLMs’ performance?
For the Design Edit and Design Repair tasks, we explore the impact of input images and codes on the results. Table VI shows the MLLM score on in Design Edit and Design Repair tasks under different context inputs. The results reveal distinct patterns regarding the utility of visual versus code information across different models and tasks.
Code-only input consistently outperforms image-only input across both tasks and all models. For Design Edit, top models (Claude-3.7, GPT-4o, Gemini-2.0) achieve highest scores with code-only input (8.40-8.43) versus image-only (7.37-7.67). Design Repair shows similar code-only superiority (6.53-6.70 vs 5.47-5.81 for top models).
However, combining code and image inputs does not yield significant improvements and occasionally results in minor performance degradation, highlighting the limitations of MLLMs in accurately localizing modification points and identifying UI issues through visual analysis.
TABLE VI: Performance under different context inputs. Both denotes combining image and code.
Model Design Edit Design Repair
Image Code Both Image Code Both
Claude-3.7 7.6764 8.4326 8.4258 5.8142 6.7014 6.8535
GPT-4o 7.3728 8.4013 8.4258 5.6968 6.5304 6.4041
Gemini-2.0 7.6430 8.4105 8.5083 5.4712 6.6726 6.2506
Llama-90B 4.9646 7.4939 6.4415 3.7860 5.3178 4.8373
Llama-11B 2.5431 6.3324 4.9111 3.6713 4.9005 3.8260
Pixtral-124B 7.6418 8.5663 8.4627 5.3816 6.5486 6.5837
Pixtral-12B 6.2005 7.6840 7.7583 4.8128 6.2159 5.9289
Qwen-72B 4.7724 8.2313 8.2442 5.3158 6.2308 6.2682
Qwen-7B 1.8098 2.9874 3.0256 1.4672 1.5225 1.3079
Finding 6: Code-only input consistently outperforms image-only input, but combining both input types does not yield much improvement. This suggests that textual code offers MLLMs with more semantic information than visual data in Design edit and Repair task.
(a) React.
(b) Vue.
(c) Angular.
Figure 7: Compilation error distribution.
VI-E RQ5: What are the limitations of MLLMs when developing framework-based webpages?
VI-E1 Compilation Errors
The compilation error indicates that MLLM does not sufficiently understand the syntax of framework-based front-end development. As illustrated in Figure 7, different MLLMs exhibit distinct error patterns across frameworks, revealing specific weaknesses in their understanding of front-end development syntax.
In React development, the top three errors are “Unexpected Token”, “Expression Expected”, and “Use Client Missing”, with Qwen-7B producing the highest number of compilation errors (95 total) while GPT-4o demonstrates superior performance with only 3 total errors.
For Vue development, “Missing End Tag”, “Unexpected EOF”, and “Attribute Error” are dominated errors. Qwen-7B shows the most compilation issues (102), while Claude-3.7 achieves near-perfect compilation success with just 1 error.
Angular development presents a different error profile, with “Incomplete Block”, “Component Import Error”, and “Component Export Error” as the primary issues. Qwen-7B continues to struggle with 73 total errors, while Pixtral-124B shows improved performance compared to other frameworks.
The error distribution reveals that MLLMs face distinct limitations when working with different frameworks: they struggle with JSX syntax parsing and React-specific expressions in React applications, encounter difficulties with template structure and attribute handling in Vue development, and show inadequate understanding of TypeScript module systems and component architecture in Angular projects. Advanced models like GPT-4o, Claude-3.7, and Pixtral-124B consistently demonstrate superior syntax understanding.
Finding 7: MLLMs exhibit framework-specific limitations: struggling with JSX parsing in React, template syntax in Vue, and TypeScript components in Angular. Advanced models demonstrate significantly better syntax comprehension across all frameworks.
To verify MLLM’s ability to solve compilation errors, we sample 30 webpages with diverse compilation errors and prompt MLLMs to fix them. The results are shown in the Table VIII. The overall average repair rate across all models and frameworks is 0.53, indicating that MLLMs still face challenges in fixing front-end errors.
TABLE VII: Compile error fix ratio of MLLMs.
Model React Vue Angular
Claude-3.7 0.70 0.40 0.70
GPT-4o 0.60 0.50 0.50
Gemini-2.0 0.60 0.80 0.80
Llama-90B 0.50 0.60 0.80
Llama-11B 0.30 0.50 0.20
Pixtral-124B 0.70 0.70 0.70
Pixtral-12B 0.50 0.60 0.40
Qwen-72B 0.70 0.50 0.50
Qwen-7B 0.20 0.10 0.20
Average 0.53 0.52 0.53
TABLE VIII: Ratio of webpages using component-based design.
Model React Vue Angular
Claude-3.7 0.23% 6% 38%
GPT-4o 0.71% 6.3% 10%
Gemini-2.0 0.7% 0.23% 41%
Llama-90B 0% 2% 5%
Llama-11B 0.48% 0.47% 1.3%
Pixtral-124B 0% 5.3% 7.7%
Pixtral-12B 0% 1.4% 2.3%
Qwen-72B 0% 17% 28%
Qwen-7B 0% 6% 40%
Average 0.24% 5% 19%
VI-E2 Component-based Implementation Limitation
Component-based design is a method of breaking down user interfaces into reusable, self-contained parts called components. This approach improves efficiency and scalability in website development. Table VIII shows the proportion of webpages implemented by MLLMs using component-based design across three frameworks. The result reveals that MLLMs demonstrate remarkably low adoption rates of component-based implementation, with average 0.24%, 5% and 19% rate on React, Vue and Angular, respectively.
Case Study. Fig. 8 shows a screenshot of the BBC website’s recently live section. Each news item follows an identical structure, making it ideal for component-based implementation with iterative rendering. However, as shown in Listing 1, the Vue code generated by MLLMs contains hardcoded, repetitive structures instead of utilizing Vue’s v-for directive. This reveals MLLMs’ insufficient understanding of component-based architecture and framework-specific syntax.
Figure 8: A webpage with multiple elements of the same structure and style.
1<!-- Item 1 -->
2<div class="border-l-2border-gray-300pl-4">
3 <div class="flex">
4 <div class="flex-1pr-4">
5 <h3 class="font-boldtext-lgmb-1">Man Utd claim huge first-leg win...</h3>
6 <p class="text-smmb-2">Manchester United take a huge step towards...</p>
7 <button class="text-xsborderborder-gray-300px-2py-1hover:bg-gray-100"> See how it played out </button>
8 </div>
9 <div class="w-1/3">
10 <img src="./placeholder.svg" alt="ManUtdvsAthleticBilbao" class="w-full" />
11 </div>
12 </div>
13</div>
14<!-- Item 2 -->
15<!-- Additional items... -->
Listing 1: Vue implementation containing repeated items.
Finding 8: MLLMs show critical deficiencies in component-based implementation and framework-specific syntax, revealing fundamental limitations in producing reusable front-end code.
VI-E3 Issue Detection Limitation
In Design Repair task, we also prompt MLLMs to judge the UI display issues described in Section IV-C. Then we calculate the issue detection accuracy by comparing the MLLMs’ outputs with the ground truth annotated by human experts. Table IX presents the UI issue detection accuracy across different MLLM models. The results demonstrate consistently poor performance of MLLMs in identifying UI issues, with average 0.2972, 0.2205 and 0.2275, 0.3403 rate on React, Vue, Angular and Vanilla, respectively.
TABLE IX: UI Issue Identification Accuracy of different MLLMs.
Model React Vue Angular Vanilla Average
Claude-3.7 0.4155 0.2654 0.3929 0.4286 0.3756
GPT-4o 0.4369 0.2870 0.4101 0.4464 0.3951
Gemini-2.0 0.4250 0.3210 0.2649 0.5417 0.3882
Llama-90B 0.2827 0.3735 0.1470 0.3571 0.2901
Llama-11B 0.0179 0.0000 0.0298 0.0119 0.0149
Pixtral-124B 0.3315 0.2747 0.3488 0.3988 0.3385
Pixtral-12B 0.3405 0.2099 0.0685 0.2804 0.2248
Qwen-72b 0.3881 0.2222 0.3851 0.3631 0.3396
Qwen-7b 0.0369 0.0309 0.0000 0.2351 0.0757
Average 0.2972 0.2205 0.2275 0.3403 0.2714
Finding 9: MLLMs struggle with identifying UI design issues accurately, with an overall average accuracy of only 0.2714 across all models and frameworks.
VII Threats to Validity
Internal validity. (1) The scores generated by MLLM-as-a-judge may have reliability concerns. To eliminate this, we carefully design the prompt and provide the model with a detailed guideline of evaluation criteria. The results of human evaluation in Section V-B also verified that the accuracy of MLLM’s score is above 90%. (2) Potential data leakage. Half of the webpages in our benchmark are sourced from closed applications (e.g., top 500 websites by Moz [37]), and the ground truth for design repair tasks is manually written by developers, minimizing data leakage risks. Moreover, the poor framework-based code generation performance, evident in syntax errors and limited use of framework-specific features, suggests that MLLM results in the benchmark are unlikely due to mere data memorization.
External validity. We only include limited frameworks of React, Vue, and Angular, due to their dominance in modern web development. These frameworks collectively represent the majority of contemporary applications and offer diverse programming paradigms: from React’s JSX to Vue’s templates and Angular’s TypeScript approach, providing comprehensive evaluation coverage.
VIII Discussion
We elicit several actionable advice from our findings for researchers and developers in this field.
For researchers: (1) Enhance framework-specific MLLM training. The poor performance on framework-specific syntax and component-based implementations underscores the need for MLLM training datasets enriched with modern web development patterns and framework-specific best practices. This would improve their practical value for applying to diverse front-end ecosystems. (2) Improve multimodal information fusion for UI tasks. Our findings reveal that for design edit and repair tasks, single code input and multimodal input perform equally well, suggesting that MLLMs currently underutilize visual information. Future work should prioritize developing more effective visual-code alignment and specialized attention mechanisms for multimodal reasoning in front-end jobs.
For developers: (1) Provide code edit location information for Design Edit and Design Repair tasks. Since MLLMs struggle with code localization, providing exact locations for edits or repairs would significantly enhance performance and reduce the cognitive burden on models to identify relevant code segments. (2) Clearly state the repair issues for repair tasks. Given MLLMs’ low accuracy in UI issue identification (27%), explicitly stating the problems allows models to focus on targeted solutions rather than inaccurate independent diagnosis. (3) Decompose complex instructions and large designs. Given that MLLM performance degrades with complex instructions and large UI mockups, practitioners can improve the practicality by breaking down complex requirements into simpler, atomic tasks and segmenting large UI designs into smaller, manageable components. This approach transforms challenging tasks into more tractable problems that align better with current MLLM capabilities.
IX Conclusion
We introduce DesignBench, the first comprehensive multi-framework multi-task benchmark for front-end code generation, encompassing React, Vue, and Angular frameworks. Beyond traditional design generation, we pioneer design edit and repair tasks. Through extensive experiments across task complexity, framework compatibility, difficulty levels, contextual factors and in-depth code-level analysis, we reveal primary limitations in current MLLMs for framework-based development and elicit several actionable advice for researchers and developers.